A process at constant T and P can be described as spontaneous if ΔG < 0 and nonspontaneous if ΔG > 0. Over what range of temperatures is the following process spontaneous? Assume that gases are at a pressure of 1 atm. (Hint: Use the data below to calculate ΔH and ΔS [assumed independent of temperature and equal to ΔH° and ΔS°, respectively] and then use the definition of ΔG.) PCl3(g) + Cl2(g) → PCl5(g) ΔHf° (kJ mol-1) -287.0 -374.9 S° (J K-1 mol-1) 311.7 223.0 364.5 ΔH° = kJ ΔS° = J K-1 The calculated value of T at which ΔG° = 0 is K. Reaction is spontaneous . Submit Answer
Answer #1

The reaction is Pelz (8) + el (g) pes (9) given aH , = -28783mal ahipels = - 374.9 kJ mo! spers = 319.757 So= 223.JK, Spelg25° = sproduet - Srunetunt. = Spels -(Spelz + sus) = 3645 Jx-(311.7 +223) JK! = 364.5Jx534.7 Jk= -170.2 Jx= -0.9702 KJK! Again AG = CH-taso At AG=0, AH Tas zo AH = Tas ...T= = AH 7100 +87.9 kJ 450 = 70.1702 KJ KAT = 516